### v2

@@ Abstract
Test-time adaptation (TTA) addresses the unforeseen distribution shifts occurring during test time. In TTA, both performance and, memory and time consumption serve as crucial considerations. A recent diffusion-based TTA approach for restoring corrupted images involves image-level updates. However, using pixel space diffusion significantly increases resource requirements compared to conventional model updating TTA approaches, revealing limitations as a TTA method. To address this, we propose a novel TTA method by leveraging a latent diffusion model (LDM) based image editing model and fine-tuning it with our newly introduced corruption modeling scheme. This scheme enhances the robustness of the diffusion model against distribution shifts by creating (clean, corrupted) image pairs and fine-tuning the model to edit corrupted images into clean ones. Moreover, we introduce a distilled variant to accelerate the model for corruption editing using only 4 network function evaluations (NFEs). We extensively validated our method across various architectures and datasets including image and video domains. Our model achieves the best performance with a 100 times faster runtime than that of a diffusion-based baseline. Furthermore, it outpaces the speed of the model updating TTA method based on data augmentation threefold, rendering an image-level updating approach more practical.
@@ 1 Introduction
To achieve efficient adaptation, we consider using Instruct-Pix2Pix (IP2P) [4], a method that utilizes a latent diffusion model (LDM) [47]. IP2P takes both text instructions and images as conditions to enable instruction-based image editing. This approach leverages not pixel space but latent space, which facilitates the efficient generation of images. However, as shown in the first row of Fig. 1, IP2P model is infeasible for TTA scenarios where out-of-domain (i.e., corrupted) images become inputs, as it can only edit in-domain images into in-domain images. Thus, to utilize IP2P model in TTA scenarios, it is crucial to enhance its robustness under distribution shifts, including test-time corruption.
@@ 2.3 Test-Time Adaptation
TTA aims to enhance inference performance with minimal resource overhead under distribution shifts. In contrast to unsupervised domain adaptation (UDA) [12, 49, 45], TTA lacks access to the source data. Moreover, unlike source-free domain adaptation [30, 28, 20], TTA has an online characteristic, obtaining target data only once through streaming. A prominent approach in TTA involves updating only a subset of parameters [59, 43, 44, 27]. However, model updating TTA methods face a risk of catastrophic forgetting [43, 61] as it lacks access to source data during training. Furthermore, the absence of clear criteria for hyperparameter selection poses a drawback, making it challenging to ensure performance in practical applications [69]. Gao et al. [13] presents diffusion-driven adaptation (DDA) to overcome these limitations. DDA utilizes a pre-trained diffusion model to transform corrupted input images into clean in-distribution images, updating the input images instead of the model. This approach enhances robustness in single-image evaluation as well as in ordered data scenarios. However, DDA falls short of meeting the efficiency requirement of TTA, as obtaining predictions for a single sample takes a long time. We greatly overcome such efficiency drawbacks. By combining LDM and CM, our model achieves higher performance and significantly reduced inference time compared to DDA. The overview of comparisons with related works are summarized in Table 1.
@@ 3.1 Diffusion Models
Diffusion models [52, 18], also known as score-based generative models [56, 57], are popular family of generative models that generate data from Gaussian noise. Specifically, these models learn to reverse a diffusion process that translates the original data distribution $p_{\text{data}}(x)$ towards a marginal distribution $q_{t}(x_{t})$, facilitated by a transition kernel defined as $q_{0t}(x_{t}|x_{0})=\mathcal{N}(x_{t}|\alpha(t)x_{0},\sigma^{2}(t)I)$, in which $\alpha(t)$ and $\sigma(t)$ are pre-defined noise schedules. Viewed from a continuous-time perspective, this diffusion process can be modeled by a stochastic differential equation (SDE) [57, 23] over the time interval $[0,T]$. To learn the reverse of SDE, diffusion models are trained to estimate score function $\epsilon_{\theta}(z_{t},t)$ with U-Net architecture [48]. Song et al. [57] show that the reverse process of SDE has its corresponding probability flow original differential equation (PF-ODE).
@@ 3.2 Consistency Models
The slow generation speed of diffusion models, due to their iterative updating nature, is a known limitation. To overcome this, the Consistency Model (CM) [55] has been introduced as a new generative model that accelerates generation to a single step or a few steps. CM operates on the principle of mapping any point from the trajectory of the PF-ODE to its destination. This is achieved through a consistency function defined as $f:(x_{t},t)\mapsto x_{\varepsilon}$, where $\varepsilon$ is a small positive value. A key aspect of CM is that the consistency function must fulfill the self-consistency property:
@@ 4 Proposed Method
A crucial consideration of Decorruptor lies in the diversity of augmentations on corruption-like augmentations during the inductive learning stage. To this end, we elucidate how we get the paired data of clean and corrupted data in 4.1, fine-tuning the models using those pairs in 4.2, distilling the model for acceleration at inference time in 4.3, and explanations on overall process in 4.4.
@@ 4.1 Corruption Modeling Scheme
Data augmentation is being used across various fields to enhance robustness against distribution shifts, including supervised [65, 64], semi-supervised [53], and self-supervised learning [5]. Similarly, in TTA, MEMO [67] applies multiple data augmentations to a single input. Notably, it utilizes $64$ data augmentations on test-time input to minimize marginal entropy and enables more stable adaptation than TENT [59].
@@
To create corrupted data, we employed prevalent data augmentation strategies on the given clean images. PIXMIX [16], for example, performs data augmentation in an on-the-fly manner based on class-agnostic complex images. Furthermore, we also consider the widely-used data augmentation method for self-supervised contrastive learning, such as SimSiam [5]. As described in Fig. 2 (a), for the crafting phase, the corrupted samples are easily crafted in a one-to-many manner from the clean image. In the following, for the learning phase, we utilized these corrupted samples for many-to-one training for corruption editing via our Decorruptor-DPM model with ‘Clean the image’ instruction $c_{T}$. Thus, our model is trained to restore mixed complex corruptions from the diverse augmented samples $c_{I}$ onto clean image $x$.
@@ 4.2.2 Scheduling Image Guidance Scale
At inference time, for Decorruptor-DPM, we utilized 20 DDIM steps and modified the image guidance scheduling to enable more effective editing. Unlike the text guidance scale, the image guidance scale is a hyper-parameter that determines how much of the input semantics are retained. Considering multi-modal conditioning guidances of the input image $c_{I}$ and text instructions $c_{T}$, CFG for our Decorruptor-DPM is as follows:
@@
It is known that if the scale is too large, the image tends to remain almost unchanged during editing, and conversely, if it is too small, the original image semantics are ignored, relying solely on text guidance for editing [4]. Since our input image is corrupted, we notice that the large guidance scale is needed at the large timestep of the pure noise phase, and a smaller guidance scale is used at the near image phase. Thus, in Eq. 7, we employed sqrt-scheduling for the image guidance scale $s_{I}$ by sampling it from $1.8$ to 00 for $t\in[T,0]$.
@@ 4.3 Decorruptor-CM: Accelerate DPM to CM
We train on the same dataset employed during the DPM training phase. While previous work [37] has only augmented PF-ODE with text guidance, we introduce augmentation with multi-modal guidance as described in Eq. 9. Multi-modal guidance has not been previously explored for diffusion distillation.
@@ 4.4 Overall Process
where $y^{pred}$ means the final prediction of our TTA method and $p_{\phi}(x)$ means the probabilistic prediction of input $x$ by the pre-trained classifier $p_{\phi}$. In contrast to Gao et al., we utilize probabilistic output after the softmax layer rather than logits for ensembling.
@@ 5 Experimental Results
In this section, we quantitatively and qualitatively validate Decorruptor. Through further analyses, we compare Decorruptor with the baseline in various aspects and demonstrate Decorruptor’s extensibility to other tasks.
@@ 5.1.1 Benchmarks
ImageNet-C [15] is a benchmark dataset generated by applying a total of 15 algorithmically generated corruptions, categorized into four groups: noise, blur, weather, and digital, to the ImageNet [10] dataset. It is utilized for evaluating the robustness under distribution shift and serves as a benchmark dataset in various tasks such as TTA and UDA. ImageNet-$\bar{\mathrm{C}}$ [39] comprises 10 corruptions that are perceptually dissimilar to those in ImageNet-C in the feature space. Additionally, to assess the effectiveness of our method in the video domain, we use UCF101-C [32], a benchmark dataset incorporating corruptions into UCF101 [58], which has a different distribution from the ImageNet.
@@ 5.1.2 Baselines
For comparison, we utilized three baselines that are not influenced by the compositions (e.g., label shifts, and batch size). DiffPure [42] is a method that employs diffusion for adversarial defense. DDA [13] uses a method that injects noise and then denoises using an in-domain pre-trained DDPM. Inspired by ILVR [6], DDA provides input that has passed through a low-pass filter as guidance and prevents catastrophic forgetting through self-ensembling. MEMO [67] is a model updating TTA approach that applies multiple data augmentations to a single input. Here, MEMO applies data augmentation to the input and tunes the classifier by minimizing the averaged prediction over augmentations. In contrast, Decorruptor uses augmentation when tuning the diffusion model for robustness against distribution shifts before adaptation.
@@ 5.1.4 Data Preparation and Model Training
For corruption crafting, we use the dataset provided by PIXMIX of fractals [41] and feature visualizations [2]. In each mixing operation in PIXMIX, we further contain SimSiam [5] transform, and various mixing sets. For model training, we initialize our model as a Stable Diffusion v1.51.51.51.5 model, and we follow the settings of IP2P for instruction fine-tuning considering image conditioning. We use ImageNet training data, with a size image of $256\times 256$. We use a total batch size of $192$ for training the model for $30,000$ steps. This training takes about $2$ days on $8$ NVIDIA A40 GPUs, and we set the universal instruction as ‘Clean the image’ for every clean-corruption pair while training. After training the DPM model, to accelerate DPM to CM, we further conduct distillation training for $24$ A40 GPU hours. Empirically, we found the distillation training for CM converges faster than training DPM.
@@ 5.2 Quantitative Evaluation
[…]
As illustrated by Tables 3 and 4, Decorruptor demonstrates the best performance across various backbone classifiers and datasets. In ImageNet-C, our 8$\times$×Decorruptor-CM consistently outperforms DDA while sustaining a shorter runtime than Decorruptor-DPM. It is noteworthy to recall that DiffPure and DDA exhibit lower performance than the source-only model on ImageNet-$\bar{\mathrm{C}}$ with Swin-T and ConvNeXt-T. It suggests that the introduction of CM enables multiple ensembles within a short time, leading to a significant improvement in performance. A detailed analysis of the ensemble is provided in the Appendix.
@@ 5.3.1 Comparisons with DDA
In Fig. 4, 4$\times$×Decorruptor-CM shows performance improvements over the source-only case in all corruptions except for pixelate and jpeg, and it outperforms DPM in most corruptions. Additionally, we employ the LPIPS [68] metric to measure the image-level perceptual similarity. As seen in Table 5, Decorruptor-DPM shows lower similarity with ImageNet-C compared to DDA, but closer similarity with ImageNet. This suggests that Decorruptor performs more edits on corrupted images compared to DDA, and the edited images becomes more cleaner images. In Fig. 5, we showcase the qualitative results of Decorruptors and DDA for a range of corruptions on ImageNet-C and ImageNet-$\bar{\mathrm{C}}$. As a result, our Decorruptor consistently outperforms DDA for all of the corruption editing.
@@ 5.3.2 Orthogonality with Model Updating Methods
We conducted experiments to explore the feasibility of combining Decorruptor with model updating methods. We compared the performance by ensembling predictions of images edited with Decorruptor-DPM to the existing model updating method. As seen in Table 6, Decorruptor demonstrated performance improvements for both TENT [59] and DeYO [27]. These outcomes suggest the potential for an advanced TTA approach through the integration of model updating and input updating methods.

### v3

@@ Abstract
Test-time adaptation (TTA) addresses the unforeseen distribution shifts occurring during test time. In TTA, performance, memory consumption, and time consumption are crucial considerations. A recent diffusion-based TTA approach for restoring corrupted images involves image-level updates. However, using pixel space diffusion significantly increases resource requirements compared to conventional model updating TTA approaches, revealing limitations as a TTA method. To address this, we propose a novel TTA method that leverages an image editing model based on a latent diffusion model (LDM) and fine-tunes it using our newly introduced corruption modeling scheme. This scheme enhances the robustness of the diffusion model against distribution shifts by creating (clean, corrupted) image pairs and fine-tuning the model to edit corrupted images into clean ones. Moreover, we introduce a distilled variant to accelerate the model for corruption editing using only 4 network function evaluations (NFEs). We extensively validated our method across various architectures and datasets including image and video domains. Our model achieves the best performance with a 100 times faster runtime than that of a diffusion-based baseline. Furthermore, it is three times faster than the previous model updating TTA method that utilizes data augmentation, making an image-level updating approach more feasible.  [^1] Project page: https://github.com/oyt9306/Decorruptor
@@ 1 Introduction
To achieve efficient adaptation, we consider using Instruct-Pix2Pix (IP2P) [5], a method that edits images in the compressed latent space. IP2P takes both text instructions and images as conditions to enable instruction-based image editing. This approach leverages not pixel space but latent space, which facilitates the efficient generation of images. However, as shown in the first row of Fig. 1, IP2P model is infeasible for TTA scenarios where out-of-domain (i.e., corrupted) images become inputs, as it can only edit in-domain images into in-domain images. Thus, to utilize IP2P model in TTA scenarios, it is crucial to enhance its robustness under distribution shifts, including test-time corruption.
@@ 2.3 Test-Time Adaptation
TTA aims to enhance inference performance with minimal resource overhead under distribution shifts. In contrast to unsupervised domain adaptation (UDA) [13, 52, 48], TTA lacks access to the source data. Moreover, unlike source-free domain adaptation [32, 30, 22], TTA has an online characteristic, obtaining target data only once through streaming. A prominent approach in TTA involves updating only a subset of parameters [62, 46, 47, 29, 45]. However, model updating TTA methods face a risk of catastrophic forgetting [46, 63] as it lacks access to source data during training. Furthermore, the absence of clear criteria for hyperparameter selection poses a drawback, making it challenging to ensure performance in practical applications [71]. Gao et al. [14] presents diffusion-driven adaptation (DDA) to overcome these limitations. DDA utilizes a pre-trained diffusion model to transform corrupted input images into clean in-distribution images, updating the input images instead of the model. This approach enhances robustness in single-image evaluation as well as in ordered data scenarios. However, DDA falls short of meeting the efficiency requirement of TTA, as obtaining predictions for a single sample takes a long time. We greatly overcome such efficiency drawbacks. By combining LDM and CM, our model achieves higher performance and significantly reduced inference time compared to DDA. The overview of comparisons with related works is summarized in Table 1.
@@ 3.1 Diffusion Models
Diffusion models [55, 20], also known as score-based generative models [59, 60], are a popular family of generative models that generate data from Gaussian noise. Specifically, these models learn to reverse a diffusion process that translates the original data distribution $p_{\text{data}}(x)$ towards a marginal distribution $q_{t}(x_{t})$, facilitated by a transition kernel defined as $q_{0t}(x_{t}|x_{0})=\mathcal{N}(x_{t}|\alpha(t)x_{0},\sigma^{2}(t)I)$, in which $\alpha(t)$ and $\sigma(t)$ are pre-defined noise schedules. Viewed from a continuous-time perspective, this diffusion process can be modeled by a stochastic differential equation (SDE) [60, 25] over the time interval $[0,T]$. To learn the reverse of SDE, diffusion models are trained to estimate score function $\epsilon_{\theta}(z_{t},t)$ with U-Net architecture [51]. Song et al. [60] show that the reverse process of SDE has its corresponding probability flow original differential equation (PF-ODE).
@@ 3.2 Consistency Models
Due to their iterative updating nature, the slow generation speed of diffusion models is a known limitation. To overcome this, the Consistency Model (CM) [58] has been introduced as a new generative model that accelerates generation to a single step or a few steps. CM operates on the principle of mapping any point from the trajectory of the PF-ODE to its destination. This is achieved through a consistency function defined as $f:(x_{t},t)\mapsto x_{\varepsilon}$, where $\varepsilon$ is a small positive value. A key aspect of CM is that the consistency function must fulfill the self-consistency property:
@@ 4 Proposed Method
A crucial consideration of Decorruptor lies in the diversity of augmentations on corruption-like augmentations during the inductive learning stage. To this end, we elucidate how we get the paired data of clean and corrupted data in 4.1, fine-tuning the models using those pairs in 4.2, distilling the model for acceleration at inference time in 4.3, and explanations on overall process in 4.4. Please refer to Appendix A.4 for the pseudo-codes of training and inference of Deccoruptor-CM.
@@ 4.1 Corruption Modeling Scheme
To create corrupted data, we employed prevalent data augmentation strategies on the given clean images. PIXMIX [18], for example, performs data augmentation in an on-the-fly manner based on class-agnostic complex images. Furthermore, we also consider the widely-used data augmentation method for self-supervised contrastive learning, such as SimSiam [6]. To augment the image, for the crafting phase, the corrupted samples are easily crafted in a one-to-many manner from the clean image. In the following, for the learning phase, we utilized these corrupted samples for many-to-one training for corruption editing via our Decorruptor-DPM model with ‘Clean the image’ instruction $c_{T}$. Thus, our model is trained to restore mixed complex corruptions from the diverse augmented samples $c_{I}$ onto clean image $x$. The connection modeling process is summarized in Fig. 2 (a).
@@ 4.2.2 Scheduling Image Guidance Scale
At inference time, for Decorruptor-DPM, we utilized 20 DDIM [57] steps and modified the image guidance scheduling to enable more effective editing. Unlike the text guidance scale, the image guidance scale is a hyper-parameter that determines how much of the input semantics are retained. Considering multi-modal conditioning guidances of the input image $c_{I}$ and text instructions $c_{T}$, CFG for our Decorruptor-DPM is as follows:
@@
It is known that if the scale is too large, the image tends to remain almost unchanged during editing, and conversely, if it is too small, the original image semantics are ignored, relying solely on text guidance for editing [5]. Since our input image is corrupted, we notice that the large guidance scale is needed at the large timestep of the pure noise phase, and a smaller guidance scale is used at the near image phase. Thus, in Eq. (7), we employed sqrt-scheduling for the image guidance scale $\omega_{I}$ by sampling it from $1.8$ to 00 for $t\in[T,0]$.
@@ 4.3 Decorruptor-CM: Accelerate DPM to CM
We train on the same dataset employed during the DPM training phase. While previous work [39] has only augmented PF-ODE with text guidance, we introduce augmentation with multi-modal guidance as described in Eq. (9).
@@ 4.4 Overall Process
where $y^{pred}$ means the final prediction of our TTA method and $p_{\phi}(x)$ means the probabilistic prediction of input $x$ by the pre-trained classifier $p_{\phi}$. In contrast to [14], we utilize probabilistic output after the softmax layer rather than logits for ensembling.
@@ 5 Experimental Results
In this section, we validate Decorruptor quantitatively and qualitatively, comparing it with the baseline and demonstrating its extensibility to other tasks.
@@ 5.1.1 Benchmarks
ImageNet-C [16] is a benchmark with 15 types of algorithmically generated corruptions in four categories: noise, blur, weather, and digital, applied to the ImageNet [11] dataset. ImageNet-$\bar{\mathrm{C}}$[41] includes 10 perceptually different corruptions. To assess our method’s effectiveness in the video domain, we use UCF101-C[34], a benchmark with corruptions applied to UCF101 [61], which has a different distribution from ImageNet.
@@ 5.1.2 Baselines
For comparison, we utilized three baselines unaffected by compositions (e.g., label shifts, batch size). DiffPure [44] employs diffusion for adversarial defense. DDA [14] uses noise injection and denoising with an in-domain pre-trained DDPM, inspired by ILVR [7], and prevents catastrophic forgetting through self-ensembling. MEMO [69] updates the model using TTA with multiple data augmentations on a single input.
@@ 5.1.4 Data Preparation and Model Training
For corruption crafting, we use the dataset provided by PIXMIX of fractals [43] and feature visualizations [2]. In each mixing operation in PIXMIX, we further apply SimSiam [6] transform, and various mixing sets. For model training, we initialize our model as a Stable Diffusion v1.51.51.51.5 model, and we follow the settings of IP2P for instruction fine-tuning considering image conditioning. We use ImageNet training data, with a size image of $256\times 256$. We use a total batch size of $192$ for training the model for $30,000$ steps. This training takes about $2$ days on $8$ NVIDIA A40 GPUs, and we set the universal instruction as ‘Clean the image’ for every clean-corruption pair while training. After training the DPM model, to accelerate DPM to CM, we further conduct distillation training for $24$ A40 GPU hours. Empirically, we found the distillation training for CM converges faster than training DPM.
@@ 5.2 Quantitative Evaluation
[…]
Tables 3 and 4 show Decorruptor’s superior performance across various classifiers and datasets. In ImageNet-C, 8$\times$×Decorruptor-CM consistently outperforms DDA with shorter runtime than Decorruptor-DPM. DiffPure and DDA perform worse than the source-only model on ImageNet-$\bar{\mathrm{C}}$ with Swin-T and ConvNeXt-T. This suggests that CM enables multiple ensembles quickly, significantly improving performance. A detailed ensemble analysis is in Appendix B.7.
@@ 5.3.1 Comparisons with DDA
We employ the LPIPS [70] metric to measure the image-level perceptual similarity. As seen in Table 5.3.3, Decorruptor-DPM shows lower similarity with ImageNet-C than DDA, but closer similarity with ImageNet. This suggests that Decorruptor performs more edits on corrupted images than DDA, and the edited images become cleaner. Moreover, in Fig. 4, we showcase the qualitative results of Decorruptors and DDA for a range of corruptions on ImageNet-C and ImageNet-$\bar{\mathrm{C}}$. As a result, our Decorruptor consistently outperforms DDA for all of the corruption editing.
@@ 5.3.2 Orthogonality with Model Updating Methods
We conducted experiments to explore the feasibility of combining Decorruptor with model updating methods. We compared the performance by ensembling predictions of images edited with Decorruptor-DPM to the existing model updating method. As seen in Table 5.3.3, Decorruptor demonstrated performance improvements for both TENT [62] and DeYO [29]. These outcomes suggest the potential for an advanced TTA approach through the integration of model updating and input updating methods. Further results are represented in the Appendix B.2.
5.3.3 Performance Trade-Off Analysis of Deccorruptor-CM
Table 5.3.3 reports the experimental results for various choices of Decorruptor-CM. When editing the input image only once for prediction, it is fast but the performance decreases by about 2-3% compared to Decorruptor-DPM. However, obtaining four edited images for each input and using their average prediction in Eq. (10) leads to higher accuracy than DPM, regardless of the architecture. Notably, even for the same corrupted image, Decorruptor allows diverse edits towards different clean images. The increasing performance gap between 1 step and 4 steps as the model size grows suggests that the quality of images is inferred to be better with 4 steps.
Table 5: LPIPS scores with clean and corrupted images.
LPIPS IN-C($\uparrow$) IN($\downarrow$)
DDA 0.421 0.608
Decorruptor-DPM 0.575 0.573
Table 6: Orthogonality with model updates.
Avg. Acc (%)
TENT 43.02
+ Decorruptor-DPM 45.52
DeYO 48.61
+ Decorruptor-DPM 49.50
Table 7: Variants of Decorruptor-CM.
Decorruptor ResNet-50 Swin-T
DPM (0.42s) 30.5 37.8
CM (1step) (0.05s) 26.8 34.5
4$\times$×CM (1step) (0.08s) 32.6 38.7
CM (4step) (0.10s) 27.5 35.6
4$\times$×CM (4step) (0.14s) 32.8 39.7
Table 8: Performance comparisons based on different source model in OOD datasets.
Source + DDA + DPM + 4$\times$×CM PIXMIX + DDA + DPM + 4$\times$×CM
VISDA-2021 acc (%) 35.7 40.2 40.9 42.0 44.0 45.4 45.6 46.1
ImageNet-A acc (%) 0.0 0.5 1.9 2.7 6.3 5.2 (-1.1) 8.1 9.8
Table 9: Performance comparisons based on different types of data augmentation methods used in the corruption modeling scheme.
Method IN-C Acc (%$\uparrow$ VISDA-2021 Acc (%$\uparrow$
PIXMIX 28.2 38.2
SimSiam 20.4 37.0
PIXMIX + SimSiam (Ours) 30.5 40.9
5.3.4 Analysis of OOD Generalization Performance
To demonstrate Decorruptor’s OOD generalization capabilities, we used the VISDA-2021 dataset [3], which includes ImageNet-C, ImageNet-R, ObjectNet, and ImageNet-A [17]. As shown in Table 8, our method outperforms both the source model and DDA on every benchmark dataset. Additionally, using Decorruptor with a single robust source model (PIXMIX) results in further performance gains, while DDA shows a performance drop on ImageNet-A. The reasons for our performance improvements in OOD are: 1) Initialization: Initializing Decorruptor with Stable Diffusion, pre-trained on the 5-billion-scale LAION dataset, enhances generalization on OOD datasets, even after fine-tuning on ImageNet. 2) Corruption Modeling Scheme: This scheme improves robustness against unseen corruption by expanding the model’s manifold by recovering corrupted images to clean images. Table 9 shows a noticeable performance difference when using PIXMIX and SimSiam, with the combination maximizing performance. This indicates that our corruption modeling scheme with effective data augmentations is beneficial for OOD generalization.
5.3.5 Multi-Modal Guidance Conditioning Analysis
We describe the benefits of using a learnable image guidance scale in CM. LCM [39] demonstrates fast convergence and significant performance improvements in distillation by using learnable guidance scales on SD, focusing on a learnable text guidance $w_{T}$. However, we found that directly implementing LCM results in undesirable outcomes for Decorruptor, which receives both text and image inputs. Recognizing the importance of conditioning the image guidance scale $w_{I}$, we proposed a new learnable multi-modal guidance $w=w_{I}\cdot w_{T}$. Detailed results are presented in Appendix C.1.
5.4 Video Test-Time Adaptation
Decorruptor demonstrates significantly improved runtime compared to DDA, making it practical for both image and video domains. We evaluated Decorruptor on a corrupted video dataset [34], applying Text2Video-Zero [28] for temporally consistent frames. Text2Video-Zero uses cross-frame attention from the first frame across the sequence for coherent editing. Detailed quantitative results are in Appendix B.4.
6 Conclusions
The existing diffusion-based image-level updating TTA approach is robust to data order and batch size variations but is impractical for real-world usage due to its slow processing speed. In response, we propose Decorruptor-DPM, leveraging a latent diffusion model for efficient memory and time utilization. Through fine-tuning via our novel corruption modeling scheme, Decorruptor-DPM possesses the capability to edit corrupted images. Additionally, we introduce Decorruptor-CM, employing consistency distillation to accelerate input updates further. Decorruptor surpasses the baseline diffusion-based approach in speed by 100 times while delivering superior performance.
6.0.1 Acknowledgement
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.2022R1A3B1077720) and the BK21 FOUR program of the Education and the Research Program for Future ICT Pioneers, Seoul National University in 2024.
Appendix 0.A Further Explanations
0.A.1 TTA with Data Augmentation Approaches
Data augmentation is being used across various fields to enhance robustness against distribution shifts, including supervised [67, 66], semi-supervised [56], and self-supervised learning [6]. Similarly, in TTA, MEMO [69] applies multiple data augmentations to a single input. Notably, it utilizes $64$ data augmentations on test-time input to minimize marginal entropy and enables more stable adaptation than TENT [62]. Here, MEMO applies data augmentation to the input and tunes the classifier by minimizing the averaged prediction over augmentations. In contrast, Decorruptor uses augmentation when training the diffusion model for robustness against distribution shifts before adaptation.
0.A.2 Detailed Contributions of Decorruptor
We clarify our Decorruptor is a classifier-agnostic generator that modifies corrupted images into clean images. Decorruptor allows for obtaining stable performance through the ensemble of multiple decorrupted images. Moreover, it can also remove corruption from images with out-of-distribution (OOD) classes, making them usable for downstream tasks. This is supported by VideoTTA results (Section 5.4) in the main text. These are key differences from other EM-based TTA approaches (e.g., EATA, SAR, and DeYO) and a single robust model (e.g., PIXMIX). Following Eq. (10), the final prediction is obtained by ensembling the predictions of the generated clean images with the original prediction. This indicates that Decorruptor can be applied orthogonally with other TTA methods that modify the original prediction. Note that the single Decorruptor checkpoint was utilized across all datasets, methods (e.g., EATA and PIXMIX), and tasks (e.g., image and video classification), demonstrating its versatility. Furthermore, Decorruptor achieves a threefold increase in speed and superior performance compared to the data augmentation-based model updating baseline.
0.A.3 Justification/Implication of Universal Prompt:
We chose the general text prompt (Clean the Image) to handle any unknown distribution shifts at test time (i.e., corruption levels and types), Other valid prompts (e.g., Decorrupt the image) will also work while they are fixed during training and inference time. To clarify it, since this prompt is fixed, our Decorruptor can be considered an image-to-image translation model that reverts corrupted images to their clean image counterparts.
0.A.4 Pseudo-codes
We provide a pseudo-code in Algorithm 1 for training Decorruptor-CM with consistency distillation [58]. Note, following LCM [39], two distinct timesteps $t_{n}$ and $t_{n+k}$, which are $k$ steps apart, are randomly selected and the same Gaussian noise $\epsilon$ are applied. The generated noisy latents $z_{t_{n}}$ and $z_{t_{n+k}}$ are represented as follows:
$z_{t_{n+k}}=\alpha(t_{n+k})z+\sigma(t_{n+k})\epsilon,\quad z_{t_{n}}=\alpha(t_{n})z+\sigma(t_{n})\epsilon.$
Note, following our multi-modal guidance scheme described in the main text, the self-consistency property can be held during distillation and the skipping-step technique can also be used. In the following, we append the inference pseudo-code of both Decorruptor-DPM and CM as described in Algorithm 2.
Algorithm 1 Decorruptor-CM Training
1:Input: Given dataset $\mathcal{D}^{(p)}$, distance metric $d(\cdot,\cdot)$, pre-trained model parameter $\theta$,learning rate $\eta$, EMA coeff $\mu$, noise schedule $\alpha(t)$, $\sigma(t)$, multi-modal guidance scale: $[w_{I,\min},w_{I,\max}]$ and $[w_{T,\min},w_{T,\max}]$, skipping interval $k$, and encoder $E(\cdot)$
2:Encode paired clean/corrupt data into the latent space: $\mathcal{D}^{(p)}_{z}=\{(z_{c1},z_{co},c)\mid z=E(x),(x_{cl},x_{co},c)\in\mathcal{D}^{(p)}\}$
[…]
4:repeat
5:     Sample $(z_{cl},z_{co},c)\sim\mathcal{D}^{(p)}_{z}$, $n\sim\mathcal{U}[1,N-k]$
6:     Sample $\epsilon\sim\mathcal{N}(0,I)$, $w_{I}$ and $w_{T}$
7:     $z_{t_{n+k}}\leftarrow\alpha(t_{n+k})z+\sigma(t_{n+k})\epsilon$
8:     $z_{t_{n}}\leftarrow\alpha(t_{n})z+\sigma(t_{n})\epsilon$
9:     Minimize Eq. (4)
10:     $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta,\theta^{-})$
11:     $\theta^{-}\leftarrow\text{stopgrad}(\mu\theta^{-}+(1-\mu)\theta)$
12:until convergence
Algorithm 2 Decorruptor-DPM/CM Inference
1:Input: Text and image guidance scales $\omega_{T}$ and $\omega_{I}$, Given text $c_{T}$ and corrupted image $c_{I}$, Noise schedule $\alpha(t)$, $\sigma(t)$, Decoder $D(\cdot)$
2:$ts$: Diffusion timesteps (20 for DPM, 4 for CM), $T$: maximum timesteps (1000)
3:$\epsilon_{\theta}$: Pre-trained DPM or CM
4:Sample $\hat{z}_{T}\sim\mathcal{N}(0;I)$, $z\leftarrow\epsilon_{\theta}(\hat{z}_{T},c_{T},c_{I},T)$
5:for $t\leftarrow ts\ldots 1$ $\triangleright$ sequence of timesteps
6:     $\hat{z}_{t}\sim\mathcal{N}(\alpha(t)z;\sigma^{2}(t)I)$
[…]
8:     $z\leftarrow\epsilon_{\theta}(\hat{z}_{t},c_{T},c_{I},t)$
9:end for
10:$\hat{x}_{0}\leftarrow D(z)$triangleright decoding latent to decorrupted image
11:return $\hat{x}_{0}$
Appendix 0.B Additional Results
0.B.1 Detailed Results of Image Corruption Editing
Tables 10 and 11 present detailed performance results of Decorruptor on ImageNet-C [16] and ImageNet-$\bar{\mathrm{C}}$ [41], respectively, using ResNet-50 [15] as the architecture. As shown in Table 10, Decorruptor demonstrates significant performance improvement for Noise and Weather corruptions. Moreover, in Fig. 5, 4$\times$×Decorruptor-CM shows performance improvements over the source-only case in all corruptions except for pixelate and jpeg, and it outperforms DPM in most corruptions.
DDA [14] also shares the limitation of not being able to properly edit some corruptions. Addressing this issue is crucial for the effectiveness of the input updating TTA [62] method. As illustrated in Table 11, Our Decorruptor shows consistent improvement for all corruptions in ImageNet-$\bar{\mathrm{C}}$. Commonly, ensembling more edited images always presents performance improvement for all corruptions in ImageNet-C and ImageNet-$\bar{\mathrm{C}}$.